A shopper pushes her 25 kg grocery cart by a force of 225 N inclined at an angle of 60° with the horizontal through a distance of 7.5 m. Find the work done by the force applied by the shopper.
1
Expert's answer
2020-12-17T09:10:47-0500

"W=Fs\cos\alpha=225\cdot7.5\cdot\cos60°=844(J)"